OpenStudy (anonymous):

Please help! Use an inverse matrix to solve each equation or system. Equation in reply.

OpenStudy (anonymous):

\[\left[\begin{matrix}4 & 1 \\ 2 & 1\end{matrix}\right]\left[\begin{matrix}x \\ y\end{matrix}\right]=\left[\begin{matrix}10\\ 6\end{matrix}\right]\]

OpenStudy (anonymous):

@pooja195 @Michele_Laino @triciaal @jigglypuff314

jigglypuff314 (jigglypuff314):

@Caseyyyy15 I can try to help you :) although I am not very good a matrices so I might make a mistake... let's first try to find the inverse matrix of \[\left[\begin{matrix}4 & 1 \\ 2 & 1\end{matrix}\right]\]do you know how to do that? :)

OpenStudy (anonymous):

is it the 1/ad-bc thing? and don't worry I'm awful too! :)

jigglypuff314 (jigglypuff314):

yeah, it's that :)

OpenStudy (anonymous):

\[\left[\begin{matrix}.5 & -.5 \\ -1 & 2\end{matrix}\right]\] is this right for the inverse?

jigglypuff314 (jigglypuff314):

yep! great job! now we multiply that by \[\left(\begin{matrix}10 \\ 6\end{matrix}\right)\]

jigglypuff314 (jigglypuff314):

do you know how to do that? :)

OpenStudy (anonymous):

on my calculator yes, but I'm needing to show work, how do you do it with the 2x1..

OpenStudy (anonymous):

i know how to do it if they were both 2x2 matrices

jigglypuff314 (jigglypuff314):

it's an idea called the "dot product" https://www.mathsisfun.com/algebra/matrix-multiplying.html \[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\left(\begin{matrix}e \\ f\end{matrix}\right) = \left(\begin{matrix}(a*e + b*f) \\ (c*e + d*f)\end{matrix}\right)\]

OpenStudy (anonymous):

okay is it.. \[\left(\begin{matrix}2 \\ 2\end{matrix}\right)\]

jigglypuff314 (jigglypuff314):

yep! that's what I got too :) (hopefully it's right) *crosses fingers* :3

OpenStudy (anonymous):

lol so is that the final answer??\[\left(\begin{matrix}x \\ y\end{matrix}\right)\] I've never seen the x and y thing so it freaked me out!

jigglypuff314 (jigglypuff314):

yeah the 2 2 is the final answer :) I used this as a guide: http://www.classzone.com/eservices/home/pdf/student/LA204EAD.pdf

OpenStudy (anonymous):

wow okay! thanks so much! :) Do you think you could help me with one more?

jigglypuff314 (jigglypuff314):

sure, I can try :)

OpenStudy (anonymous):

 is how it looks on my paper and its the same instructions

jigglypuff314 (jigglypuff314):

then you would get the matrix problem: \[\left[\begin{matrix}1 & 2\\ 2& 4\end{matrix}\right]\left[\begin{matrix}x\\ y\end{matrix}\right]=\left[\begin{matrix}15 \\ 30 \end{matrix}\right]\]

jigglypuff314 (jigglypuff314):

*blinks* oh wait

jigglypuff314 (jigglypuff314):

the 1/(ad - bc) would get you 1/0 that would mean that they are the same equation aka infinitely many solutions

OpenStudy (anonymous):

i thought it made since until you said that. I'm confused now..

OpenStudy (anonymous):

what are you getting the inverse of to start?

jigglypuff314 (jigglypuff314):

try doing the 1/(ad - bc) I got one divided by zero from just that but it is impossible to divide by zero

OpenStudy (anonymous):

oh from the equation you got okay one sec

OpenStudy (anonymous):

i got zero too! so is it no solution?

jigglypuff314 (jigglypuff314):

it's infinitely many solutions because they are the same equation

OpenStudy (anonymous):

interesting okay thank you so much! You rock!

jigglypuff314 (jigglypuff314):

I'm glad I could help ^_^